# Rolling with the Punches: Resilient Contrastive Pre-training under Non-Stationary Drift

### v2

@@ 2.2 Causal Interventional Contrastive Objective
$P_{t}(Y|\text{do}(X))=\sum_{i}^{|D|}P_{t}(Y|X,B=h(X,d_{i}))P_{t}(d_{i})$ (5)
In Eq. (5), $P_{t}(Y|\text{do}(X))$ compels $X$ to equally consider each confounding factor $d_{i}$ and integrate them collectively to predict $Y$. Through this approach, the original classifier $P_{t}(Y|X)$ is replaced by $P_{t}(Y|\text{do}(X))$ at the iteration $t$, as shown:
$P_{t}(Y|\text{do}(X))\triangleq P_{t}(Y|X)$ (6)
The de-confound classifier mitigates the confounding influence and captures the genuine causality from $X$ to $Y$, thereby enhancing the quality of the contrastive pre-training. Nevertheless, Eq. (5) necessitates costly sampling to approximate $P(Y|\text{do}(X))$ when implementing it in contrastive pre-training, resulting in impractical training times. Fortunately, the Normalized Weighted Geometric Mean (NWGM) (xuShowAttendTell2016; liuShowDeconfoundTell2022; yangDeconfoundedImageCaptioning2023) explores an alternative to approximate Eq. (5) in a single feed-forward process as following:
$P_{t}(Y|\text{do}(X))\approx P_{t}(Y|X,B=\sum_{i}^{|D|}h(X,d_{i})P_{t}(d_{i}))$ (7)
$L(g(\tilde{x}^{i}),m(\hat{x}^{i}))=P_{t}(Y|X,\sum_{i}^{|D|}h(X,d_{i})P_{t}(d_{i}))$ (8)
$=P_{t}(Y|g(\tilde{x}^{i}),\text{Softmax}(h(g(\tilde{x}^{i})m(\hat{x}^{i}))))$
2.3 Resilient Contrastive Pre-training
Upon achieving the causal interventional objective for contrastive learning, we devise a resilient contrastive pre-training approach to accommodate the data stream of concept drift, as depicted in Fig. 3.
Figure 3: The workflow of our causal contrastive pre-training under concept drift streaming. Within the data streaming, a large batch size is opted for a wider drift adaptation window sliding to adapt changes in data distribution. Undergoes various random augmentations, the transformed instances from the identical sample are feature-extracted by both the encoder and the momentum encoder to get the key and value, respectively. An MLP head is utilized to obtain the query of the encoder features. Subsequently, causal intervention is utilized to alleviate concept drift in the data stream within the adaptation window, resulting in the acquisition of two objects for contrastive learning.
Following the student-teacher architecture of MoCo v3 (chenEmpiricalStudyTraining2021), an encoder and a momentum encoder are utilized for causal interventional contrastive pre-training. Within the drift adaption window $[t,t+\tau]$ of the data stream, two distinct crops of each image are utilized under random data augmentation to produce two contrasting views for comparison. The encoder extracts the image feature $v=g(\tilde{x}^{i})$, while the momentum encoder is leveraged for the stability of training as the teacher network, generating the features $k=m(\hat{x}^{i})$. The momentum encoder has the same structure as the encoder, driven by the exponential moving average (EMA) as Eq.(2). Obtained the image feature $v$, an MLP head produce the projection of the encoder feature as $q=h(g(\tilde{x}^{i}))$. Consequently, the intervention module is employed to eliminate latent biases based on features $q$, $v$ and $k$ within the concept drift data stream, and create the contrastive objective $c=\text{Softmax}(q\cdot k)\cdot v$ as Eq. (8). In particular, we use cross-fusion to estimate the concept drift in the data stream, i.e. $q$ and $k$ are obtained from different views, helping the stability of the pre-training. Finally, the InfoNCE loss (oordRepresentationLearningContrastive2019) is adopted to drive the whole network as follows:
$\mathcal{L}=-\log\frac{\exp{(c_{1}\cdot c_{2}^{i+}/\tau)}}{\sum_{i=t}^{t+\tau}\exp{(c_{1}\cdot c_{2}^{i}/\tau)}}$ (9)
where $\tau$ is a temperature hyper-parameter (wuUnsupervisedFeatureLearning2018a). The sum is over one positive and all negative samples within the drift adaptation window $[t,t+\tau]$ of the data stream.
Intuitively, $q$ akin to a ”query” (heMomentumContrastUnsupervised2020), the momentum output $k$ acts as the ”key” for sampling key drift of the data stream, and the ”value” $v$ resembles a blend of the essential contrastive object and the concept drift. Therefore, the goal of causal intervention is to mitigate the bias of the contrastive objective through the drift sampling within the adaption window. We find the derived causal intervention module has a similar structure to the self-attention mechanism (vaswaniAttentionAllYou2017). Thus, it is argued that the causal intervention contrastive learning has the ability to capture long-sequence relationships similar to self-attention, which aids in capturing concept drift within a broader drift adaptation window in contrastive pre-training. However, it is important to highlight that we employ ”key” and ”query” sampling for concept drift within the data stream to mitigate bias in the ”value”, which is different from the aim of sequence position modelling in the self-attention mechanism.
2.4 Simple and Scalable Implementation
Our implementation of the proposed causal interventional contrastive pre-training is highly efficient and simple, requiring minimal specialized operations. As illustrated in the workflow depicted in Fig. 3, we have made only minor adjustments to the MoCo v3 code (chenEmpiricalStudyTraining2021). These modifications primarily pertain to the interventional steps following the acquisition of image feature embeddings from the encoder and the momentum encoder. Additionally, we have incorporated the masked contrastive strategy proposed in (yangMaskedImageContrastive2024) to facilitate a large batch size which enables the acquisition of the border drift adaption window.
Table 1: Evaluation results of fine-tuning on long-tailed classification tasks with ImageNet-LT (liuLargeScaleLongTailedRecognition2019) and iNatualist2018 (vanhornINaturalistSpeciesClassification2018a). The best-performing contrastive pre-training results are highlighted in red. Many, Medium and Few denote the evaluated splits of many-shot ($>$100 training samples), medium-shot (20-100 samples) and few-shot ($<$20 samples). Top-1 accuracy is applied to evaluate the performance of different methods. $\dagger$ denotes methods that are adjusted for the long-tailed fine-tuning.
ImageNet-LT iNaturalist2018
Methods Backbones Many Medium Few All Many Medium Few All
Training from scratch
cRT (kangDecouplingRepresentationClassifier2019) 61.8 46.2 27.3 49.6 69.0 66.0 63.2 65.2
LWS (kangDecouplingRepresentationClassifier2019) 60.2 47.2 30.3 49.9 65.0 66.3 65.5 65.9
RIDE (wangLongtailedRecognitionRouting2020) 68.2 53.8 36.0 56.9 70.9 72.4 73.1 72.6
PaCo (cuiParametricContrastiveLearning2021) ResNet-50 68.2 58.7 41.0 60.0 70.3 73.2 73.6 73.2
Masked Image Modeling Pre-training
MAE (heMaskedAutoencodersAre2022) 74.7 48.2 19.4 54.5 79.6 70.8 65.0 69.4
LiVT $\dagger$ (xuLearningImbalancedData2023) ViT-B/16 73.6 56.4 41.0 60.9 78.9 76.5 74.8 76.1
Contrastive Learning Pre-training
ViT (dosovitskiyImageWorth16x162021) 50.5 23.5 6.9 31.6 65.4 55.3 50.9 54.6
DeiT (touvronDeiTIIIRevenge2022) 70.4 40.9 12.8 48.4 72.9 62.8 55.8 61.0
BYOL (grillBootstrapYourOwn2020) 30.6 8.6 0.9 15.9 28.4 22.6 24.5 24.0
DINO (caronEmergingPropertiesSelfSupervised2021) 64.8 35.5 11.3 43.3 76.7 67.1 61.2 65.5
DINO v2 (zhouDINOWMWorldModels2024) 68.4 36.9 11.7 45.4 71.9 62.7 58.2 61.6
MoCo v3 (chenEmpiricalStudyTraining2021) ViT-B/16 70.8 40.7 14.3 48.5 76.6 65.8 62.8 65.6
Ours 72.2 43.1 16.0 50.3 77.7 68.7 63.8 67.5
@@ 2.4 Simple and Scalable Implementation
ImageNet-LT (liuLargeScaleLongTailedRecognition2019)
BYOL (grillBootstrapYourOwn2020) 8.5 0.5 0.1 3.5
DINO (caronEmergingPropertiesSelfSupervised2021) 47.4 20.5 0.1 27.9
DINO v2 (zhouDINOWMWorldModels2024) 49.4 23.4 5.7 30.8
MoCo v3 (chenEmpiricalStudyTraining2021) 45.8 18.7 0.0 27.1
Ours 52.7 24.0 7.9 32.8
iNaturalist2018 (vanhornINaturalistSpeciesClassification2018a)
BYOL (grillBootstrapYourOwn2020) 7.0 3.1 2.5 3.2
DINO (caronEmergingPropertiesSelfSupervised2021) 43.1 33.5 32.3 34.0
DINO v2 (zhouDINOWMWorldModels2024) 36.3 30.2 30.0 30.7
MoCo v3 (chenEmpiricalStudyTraining2021) 32.9 25.4 22.4 24.9
Ours 48.4 37.0 35.1 37.3
Furthermore, in comparison to prior approaches, causal contrastive learning presents exceptional scalability. Its scalability is twofold: firstly, it accommodates the inclusion of numerous drift datasets in the pre-training. Following the scaling law (muennighoffScalingDataConstrainedLanguage2023), augmenting data correlates with an enhancement in model performance. Causal contrastive pre-training guarantees that model performance is bolstered by increasing data without succumbing to the effects of concept drift. Secondly, it strengthens the stability of larger models during drift pre-training, enabling the pre-training of larger models with more parameters and deeper layers.
@@ 3 Experiments
In this section, we initially showcase the robust performance of our resilient contrastive pre-training in the downstream task of long-tailed classification under drift pre-training, with results of fine-tuning and linear probing. Subsequently, from the perspective of generalization, we conduct experiments of out-of-distribution (OOD) detection and domain shift, to assess the resilience of the model against gradual drift and sudden drift. Following that, the feature embedding space of pre-trained models is intuitively visualized with degrees, illustrating how resilient contrastive pre-training mitigates the concept drift within the feature space. Finally, we present the scaling ability of our method facing with the drift pre-training. More detailed experimental implementations are given in Appendix B.2.
@@ 3.1 Navigating Tailed Drift Pre-training
We compare our proposed resilient contrastive pre-training method with other models to explicitly demonstrate its superior performance in tailed drift pre-training. Two large-scale datasets, namely the ImageNet-LT (liuLargeScaleLongTailedRecognition2019) and iNaturalist 2018 (vanhornINaturalistSpeciesClassification2018a) are utilized as source datasets to perform pre-training, fine-tuning and linear probing, respectively. To effectively illustrate the efficacy of our resilient contrastive pre-training in mitigating tail drift, we follow the criterion and metrics of long-tailed classification (yangTdistributedSphericalFeature2023) to report the Top-1 accuracy across various splits, including Many split with over 100 training samples, Medium split with 20-100 training samples, and Few split with fewer than 20 training samples.
@@
In terms of fine-tuning results shown in Table 1, we compare mainstream contrastive pre-training methods, namely DeiT (touvronDeiTIIIRevenge2022), BYOL (grillBootstrapYourOwn2020) , DINO (caronEmergingPropertiesSelfSupervised2021) , DINO v2 (zhouDINOWMWorldModels2024), MoCo v3 (chenEmpiricalStudyTraining2021) with the baseline of ViT (dosovitskiyImageWorth16x162021), and our resilient contrastive learning exhibits superior results beyond other contrastive pre-training methods. Compared to other methods that employ the student-teacher paradigm, such as BYOL, DINO, DINO v2, and MoCo v3, our resilient contrastive pre-training method demonstrates additional performance enhancements on both the Medium and Few splits, especially in the iNaturalist2018 dataset. It is demonstrated that the proposed causal interventional objective significantly alleviates the accumulated bias due to momentum updates of the teacher network. Meanwhile, under the concept drift scenario, BYOL without negative sample pairs has the inferior performance of long-tailed classification to the baseline ViT, while it is not noticeable in equilibrium (chenEmpiricalStudyTraining2021). We argue that, in drift pre-training, the essence of negative sample pairs is similar to the causal interventional objective, taking into account the causal relationship of all drifts for each positive sample like de-confounding operation in Eq. (5) to delineate the boundary of the feature space. Relying solely on maximizing positive agreement as a pre-training strategy will cause the model to be overwhelmed by the head category, eventually exacerbating concept drift.
@@
Besides, we also provide other fine-tuning results of long-tailed classification under scratch training and masked image modelling pre-training for comprehensive analysis, as shown in Table 1. It is worth noting that our focus is on the impact of drift environments on contrastive pre-training. Therefore, we exclusively utilized resilient contrastive pre-training solely during the pre-training phase, without making any adjustments to the downstream classification tasks. Consequently, in comparison to other methods, all contrastive pre-training approaches exhibit relatively lower performance. In the context of the masked image modeling strategy, it is found that contrastive learning is more susceptible to concept drift in the data stream. we attribute it to the fact that contrastive learning relies more heavily on global information within the data flow to construct feature representations. Encountering concept drift within the data stream leads to accumulated deviations in momentum updates of contrastive learning, consequently resulting in degraded performance.
@@
Table 3: Evaluation results of domain shift on ImageNet-V2 (rechtImageNetClassifiersGeneralize2019), ImageNet-Sketch (wangLearningRobustGlobal2019) and ImageNet-R (hendrycksManyFacesRobustness2021). The best-performing models are highlighted in red. Many, Medium and Few denote the evaluated splits of many-shot ($>$100 training samples), medium-shot (20-100 samples) and few-shot ($<$20 samples). Top-1 and Top-5 accuracy are applied to evaluate the performance of different methods. We compare our methods with other contrastive pre-training models, including BYOL (grillBootstrapYourOwn2020), DINO (caronEmergingPropertiesSelfSupervised2021), DINO v2 (zhouDINOWMWorldModels2024) and MoCo v3 (chenEmpiricalStudyTraining2021).
Methods Many Medium Few Top-1 Top-5
ImageNet-V2 (rechtImageNetClassifiersGeneralize2019)
BYOL 29.7 7.8 1.0 15.2 31.8
DINO 63.4 33.6 11.6 41.9 65.1
DINO v2 66.6 35.4 11.9 44.0 66.9
MoCo v3 62.9 37.3 12.6 43.1 66.0
Ours 69.7 41.2 14.9 48.3 72.1
ImageNet-Sketch (wangLearningRobustGlobal2019)
BYOL 1.9 0.2 0.0 0.8 2.8
DINO 12.4 5.0 0.5 7.1 16.2
DINO v2 17.1 6.8 1.4 9.9 20.8
MoCo v3 17.6 7.6 1.4 10.4 21.9
Ours 22.2 9.7 2.3 13.4 27.0
ImageNet-R (hendrycksManyFacesRobustness2021)
BYOL 4.2 1.7 0.0 2.7 6.8
DINO 13.4 7.6 0.4 9.5 18.3
DINO v2 15.8 8.6 0.7 11.1 21.1
MoCo v3 17.2 9.8 1.0 12.3 22.8
Ours 21.3 12.5 1.6 15.5 27.5
3.2 Taming OOD Drift and Domain Drift in Downstream Tasks

### v3

@@ 2.3 Resilient Contrastive Pre-training
ImageNet-LT [liuLargeScaleLongTailedRecognition2019]
BYOL [grillBootstrapYourOwn2020] 8.5 0.5 0.1 3.5
DINO [caronEmergingPropertiesSelfSupervised2021] 47.4 20.5 0.1 27.9
DINO v2 [zhouDINOWMWorldModels2024] 49.4 23.4 5.7 30.8
MoCo v3 [chenEmpiricalStudyTraining2021] 45.8 18.7 0.0 27.1
ReLIC [mitrovicrepresentation] 3.2 21.0 47.1 28.6
[…]
iNaturalist2018 [vanhornINaturalistSpeciesClassification2018a]
BYOL [grillBootstrapYourOwn2020] 7.0 3.1 2.5 3.2
DINO [caronEmergingPropertiesSelfSupervised2021] 43.1 33.5 32.3 34.0
DINO v2 [zhouDINOWMWorldModels2024] 36.3 30.2 30.0 30.7
MoCo v3 [chenEmpiricalStudyTraining2021] 32.9 25.4 22.4 24.9
[…]
To realize the causal interventional objective for CDP, we introduce Resilient Contrastive Pre-training (RCP), an approach tailored for learning from data streams exhibiting concept drift. As depicted in Fig. 3, RCP adapts the student-teacher architecture of MoCo v3 [chenEmpiricalStudyTraining2021], utilizing an encoder ($g$) and a momentum encoder ($m$). Within each drift adaptation window $t,t+\tau]$, two augmented views ($\tilde{x}$, $\hat{x}$) are generated per image. The student encoder $g$ processes one view into features v=g(x~)v=g($\tilde${x}), which an MLP head then projects into a query $q=h(v)$. The momentum encoder $m$, structurally identical to $g$ and updated via EMA (Eq. (2)), processes the other view to produce a key k=m(x^)k=m($\hat${x}), ensuring stable targets and enhancing training stability. Central to RCP is an intervention module that processes $q$, $v$, and $k$ to mitigate latent biases from concept drift. This module computes an interventional score component $c=\text{Softmax}(q\cdot k)\cdot v$, which embodies aspects of our causal objective (Eq. (8)). The overall network is then trained using an InfoNCE loss [oordRepresentationLearningContrastive2019]:
$\mathcal{L}=-\log\frac{\exp{(c_{1}\cdot c_{2}^{i+}/\tau)}}{\sum_{i=t}^{t+\tau}\exp{(c_{1}\cdot c_{2}^{i}/\tau)}},$ (9)
where $\tau$ is a temperature hyper-parameter [wuUnsupervisedFeatureLearning2018a]. The sum is over one positive and all negative samples within the drift adaptation window $t,t+\tau]$ of the data stream.
Intuitively, $q$ akin to a ”query” [heMomentumContrastUnsupervised2020], the momentum output $k$ acts as the ”key” for sampling key drift of the data stream, and the ”value” $v$ resembles a blend of the essential contrastive object and the concept drift. Therefore, the goal of causal intervention is to mitigate the bias of the contrastive objective through the drift sampling within the adaption window. We find the derived causal intervention module has a similar structure to the self-attention mechanism [vaswaniAttentionAllYou2017]. Thus, this query-key-value mechanism structurally resembles self-attention [vaswaniAttentionAllYou2017], suggesting an inherent capability to model contextual dependencies useful for tracking drift. However, it is important to highlight that we employ ”key” and ”query” sampling for concept drift within the data stream to mitigate bias in the ”value”, which is different from the aim of sequence position modelling in the self-attention mechanism. Furthermore, our RCP method is simple to implement, efficient and scalable, with more details provided in Appendix.
@@ 3 Experiments
In this section, we initially showcase the robust performance of our RCP in the downstream task of long-tailed classification under drift pre-training, with results of fine-tuning and linear probing. Subsequently, from the perspective of generalization, we conduct experiments of out-of-distribution (OOD) detection and domain shift, to assess the resilience of the model against gradual drift and sudden drift. Following that, the feature embedding space of pre-trained models is intuitively visualized with degrees, illustrating how RCP mitigates the concept drift within the feature space. Finally, we present the scaling ability of our method facing with the drift pre-training.
It is worth noting that we do not assume any type of drift. We chose several benchmarks with known drifts, such as long tail, OOD, and domain shift, only for comparison with existing benchmarks, and visually illustrated the effectiveness of our model. However, we did not optimize for specific drift types during training. More detailed experimental implementations are given in Appendix.
@@ 3.1 Navigating Tailed Drift Pre-training
We compare our proposed RCP with other models to explicitly demonstrate its superior performance in tailed drift pre-training. Two large-scale datasets, ImageNet-LT [liuLargeScaleLongTailedRecognition2019] and iNaturalist 2018 [vanhornINaturalistSpeciesClassification2018a], are utilized as source datasets to perform pre-training, fine-tuning and linear probing, respectively. To effectively illustrate the efficacy of RCP in mitigating tail drift, we follow the criterion and metrics of long-tailed classification [yangTdistributedSphericalFeature2023] to report the Top-1 accuracy across various splits, including Many split with over 100 training samples, Medium split with 20-100 training samples, and Few split with fewer than 20 training samples.
@@
In terms of fine-tuning results shown in Table 1, we compare mainstream contrastive pre-training methods, namely DeiT [touvronDeiTIIIRevenge2022], BYOL [grillBootstrapYourOwn2020] , DINO [caronEmergingPropertiesSelfSupervised2021] , DINO v2 [zhouDINOWMWorldModels2024], MoCo v3 [chenEmpiricalStudyTraining2021] with the baseline of ViT [dosovitskiyImageWorth16x162021], and our RCP exhibits superior results beyond other contrastive pre-training methods. Compared to other methods that employ the student-teacher paradigm, such as BYOL, DINO, DINO v2, and MoCo v3, our RCP method demonstrates additional performance enhancements on both the Medium and Few splits, especially in the iNaturalist2018 dataset. It is demonstrated that the proposed causal interventional objective significantly alleviates the accumulated bias due to momentum updates of the teacher network. Meanwhile, under the concept drift scenario, BYOL without negative sample pairs has the inferior performance of long-tailed classification to the baseline ViT, while it is not noticeable in equilibrium [chenEmpiricalStudyTraining2021]. We argue that, in drift pre-training, the essence of negative sample pairs is similar to the causal interventional objective, taking into account the causal relationship of all drifts for each positive sample like de-confounding operation in Eq. (5) to delineate the boundary of the feature space. Relying solely on maximizing positive agreement as a pre-training strategy will cause the model to be overwhelmed by the head category, eventually exacerbating concept drift.
@@
Besides, we also provide other fine-tuning results of long-tailed classification under scratch training and masked image modelling pre-training for comprehensive analysis, as shown in Table 1. It is worth noting that our focus is on the impact of drift environments on contrastive pre-training. Therefore, we exclusively utilized RCP solely during the pre-training phase, without making any adjustments to the downstream classification tasks. Consequently, in comparison to other methods, all contrastive pre-training approaches exhibit relatively lower performance. In the context of the masked image modeling strategy, it is found that contrastive learning is more susceptible to concept drift in the data stream. We attribute it to the fact that contrastive learning relies more heavily on global information within the data flow to construct feature representations.
Table 3: Evaluation results of domain shift on ImageNet-V2 [rechtImageNetClassifiersGeneralize2019], ImageNet-Sketch [wangLearningRobustGlobal2019] and ImageNet-R [hendrycksManyFacesRobustness2021]. The best-performing models are highlighted in red. Top-1/-5 accuracy are applied to evaluate the performance.
[…]
@@
3.2 Taming ID and OOD Drift
To validate the generalization capability of RCP under drifting scenarios, we conducted experiments from the perspectives of domain shift and out-of-distribution (OOD) detection.
In terms of domain shift generalization as exhibited in Table 4, our primary focus is to assess the extent of the drift environment on the feature representation constructed during pre-training. Therefore, we select three subsets of ImageNet [russakovskyImagenetLargeScale2015] to validate the domain shift performance, namely, ImageNet-V2 [rechtImageNetClassifiersGeneralize2019], ImageNet-Sketch [wangLearningRobustGlobal2019] and ImageNet-R [hendrycksManyFacesRobustness2021]. Our results achieve superior results across three datasets, investigating the effectiveness of the proposed RCP method. It corroborates that, the feature space we construct in pre-training abstracts and captures the essential information of images, while disregarding any unexpected interference from the concept drift within the data stream. Moreover, judging from the results of Few split in ImageNet-Sketch and ImageNet-R, the majority of contrastive pre-training methods struggle to differentiate the tail categories. This difficulty arises from two main challenges. First, the bias induced by tail drift tends to skew models toward favoring head categories. On the flip side, the small sample size of tail categories limits the ability of pre-training methods to accurately extract their fundamental features, a problem that is exacerbated in the presence of domain shifts. Accordingly, our outstanding results demonstrate our proficiency in mitigating tail drift and extracting crucial features from limited samples.
Concerning out-of-distribution (OOD) detection, we also evaluate the capability of our RCP approach to delineate the boundaries of the feature space, as presented in Table 4. While exhibiting slightly lower performance than DINO on ImageNet-O, we significantly outperform contrastive pre-training methods on the other three out-of-distribution datasets, and surpasses other approaches in overall performance. It demonstrates that our model possesses the benefits of intra-class compactness and inter-class separability in feature representation. Meanwhile, it is worth noting that in larger-scale balanced pre-training datasets, such as ImageNet-21k [ridnikImageNet21KPretrainingMasses2021], our approach outperforms numerous traditional OOD detection methods in overall performance, such as ViM [wang2022vim]. It underscores our RCP method can effectively alleviate concept drift bias in the model while also enhancing the characterization of the model’s decision boundary.
Table 5: Evaluation results of different contrastive learning strategies in the stage of pre-training from three perspectives: ID intra-class compactness, ID inter-class separability and the separability between ID and OOD categories. The cosine metric is utilized to measure these distances, which is expressed as average degrees. ImageNet-LT is utilized for pre-training. Best results are highlighted in red.
ID Intra Compactness $\downarrow$ ID Inter Separability $\uparrow$ ID vs. OOD Separability $\uparrow$
Pre-training Many Medium Few All Many Medium Few All Many Medium Few All
BYOL [grillBootstrapYourOwn2020] 53.1 59.1 59.9 56.9 81.5 79.2 78.3 80.0 75.7 71.9 70.3 73.1
DINO [caronEmergingPropertiesSelfSupervised2021] 34.2 51.4 59.0 45.9 89.0 88.4 87.6 88.5 86.2 84.4 81.4 84.7
DINO v2 [zhouDINOWMWorldModels2024] 31.2 50.1 57.8 43.9 88.9 88.1 87.2 88.3 86.0 83.7 80.5 84.1
MoCo v3 [chenEmpiricalStudyTraining2021] 29.1 47.8 57.8 42.8 89.3 88.8 88.2 88.9 86.9 85.7 83.0 85.8